OpenStudy (anonymous):

what will be area of triangle whose points are (0,0) (0,h) and (h,0)

OpenStudy (anonymous):

\[h ^{2}/2\]

OpenStudy (anonymous):

(h^2)/2

OpenStudy (bahrom7893):

plot it it will be a right triangle, so the area is (1/2)*base*height = h*h*(1/2) = h^2/2

OpenStudy (bahrom7893):

fan us all if we helped, thanks! =)